# VELMA: Verbalization Embodiment of LLM Agents for Vision and Language Navigation in Street View

### v1

@@ 6.2 Few-Shot Learning Results
Figure 5: Key point accuracy (KPA) for 2-shot in-context learning of large language models with increasing parameter count. The ${\tt{FORWARD}}$-Baseline predicts only walking forward and is better than random actions.
The proposed text-only interface allows us to use large language models as reasoners without updating their weights or fusing image representations. The prompt consists of a short task description and two in-context examples (2-shot). The examples are full text sequences for randomly selected navigation instances in the training set. The plots in Figure 5 show that performance scales with parameter count as expected. The ${\tt{FORWARD}}$-baseline reveals that OPT (Zhang et al. 2022) can barely compete with this baseline yielding nonsensical predictions, even at model sizes of 65 billion parameters. The LLaMa model family on the other hand shows substantial navigation skills reaching 45.7 and 48.7 key point accuracy (KPA) on Touchdown and Map2seq, respectively. However, this KPA score only translates to task completion (TC) rates of 1.6 and 2.5, revealing that the model is not able to consistently predict correct actions throughout the whole navigation trajectory. The only few-shot LLMs that achieve substantial TC rates are GPT-3 and GPT-4 (OpenAI 2023). As listed in Table 2, VELMA-GPT-4 achieves the best results for the 2-shot setting. It reaches 44% and 77% of the TC rate reported for the previous state-of-the-art model ORAR* (Schumann and Riezler 2022) which is a seq-to-seq model that has direct access to image features and was trained on the full training set. In contrast, the GPT models act as a blind agent that solely relies on observation descriptions produced by the verbalizer. This is remarkable because LLMs are not explicitly trained to experience embodiment in a visual environment. This is emergent behavior unearthed by verbalizing the VLN task. We also observe that the GPT models use the ${\tt{TURN\_AROUND}}$ action in appropriate ways, even when the in-context examples do not demonstrate the usage of it. This emphasizes the effectiveness of intuitive communication with the environment.
@@ 6.3 Finetuning Results
To further explore the capabilities of the proposed LLM agent, we finetune LLaMa-7b on all training instances of the respective dataset, denoted by VELMA-FT in Table 2. Each training instance is the full text sequence after following the gold path. The visibility of landmarks is determined by the landmark scorer during training because gold annotations are not available. There are 6,770 training instances for Touchdown and 5,737 for Map2seq. We finetune for 20 epochs using LoRA (Hu et al. 2022) to adapt attention query and value projections. The best model is selected by task completion on the development set. The resulting agent outperforms the previous state-of-the-art model ORAR* by 10% and 16% relative TC rate. Comparing ORAR* which fuses image features at the vector level to VELMA-FT which finetunes on verbalizations of observations, shows that the text-based environment observations are less prone to overfitting.
@@
A navigation task is successfully completed if the agent stops at either the goal location or an adjacent neighboring node. Training the agent with teacher-forcing to exactly follow the gold route penalizes the agent for stopping one step short or one step past the target node, despite accomplishing the navigation objective. Furthermore, the agent can not learn to recover from incorrect decisions during inference. We thus train the agent to directly optimize the TC metric while also feeding it its own actions during training, called VELMA-RBL in Table 2. The procedure for VELMA-RBL is inspired by response-based learning (Clarke et al. 2010) and imitation learning (Ross, Gordon, and Bagnell 2011) and is outlined in Algorithm 1. The loss for an instance at training step $j$ is either computed by teacher forcing the gold action sequence $\hat{\mathbf{a}}$, or by student forcing, determined by a mixing parameter $\lambda$. In the latter case, the agent freely moves around, executing the actions decoded using its current model weights $\theta_{j}$. If the agent successfully stops within one neighboring node of the target location, the predicted action sequence $\mathbf{a}_{j}$ is considered correct and used as the reference to train the agent. If the agent stops at the wrong location, an oracle path is computed to provide the optimal counterfactual action at each step in the trajectory. In our case, the oracle’s optimal next action is computed as the shortest path to the goal location. We set $\lambda=0.5$ to collect training losses in a batch evenly from both training strategies. Manually inspecting trajectories produced by the trained agent, we found improvements of following instructions that have stopping criteria like ”Stop a few steps before Y.” or ”Stop at X. If you see Y you have gone too far.”. In both cases, the agent learned to walk past the uncertain stopping location and to invoke the ${\tt{TURN\_AROUND}}$ action in order to walk back once landmark Y appeared. Overall the task completion rate increases significantly using this training procedure with 15% and 11% relative improvements for Touchdown and Map2seq respectively. Overall, our contributions in this work amount to an absolute increase of task completion by 9.2 and 11.8 over the previously reported state-of-the-art.
@@ 6.4 Image Ablation
Touchdown Map2seq
Image Model SPD↓ KPA↑ TC↑ SPD↓ KPA↑ TC↑
no image 27.4 45.5 14.7 9.7 76.4 35.2
CLIP 21.3 57.4 19.5 9.8 76.9 37.2
OpenCLIP 18.6 61.3 22.6 9.8 76.8 38.2
@@ 7 Conclusion
We introduced VELMA, an agent for urban vision and language navigation, which utilizes a large language model to infer its next action. The LLM is continuously queried with a text prompt that verbalizes the task description, navigation instructions, visual observations, and past trajectory of the agent. In order to include observed landmarks in the prompt, we propose an unsupervised pipeline that extracts landmarks from the instructions and determines their visibility in the current panorama view based on thresholded CLIP scores. We evaluate the embodied LLM agent in a modified version of the commonly used Touchdown environment based on Street View. One proposed modification is fixing a problem at intersections that led to incorrect alignments of action sequences, and another modification adds the ${\tt{TURN\_AROUND}}$ action which provides a more intuitive way to communicate with the environment. The proposed agent achieves strong few-shot in-context learning results of 10 and 23 task completion rates for Touchdown and Map2seq, respectively, and yields new state-of-the-art results of 24 and 42 task completion rates when finetuned on the full training set.
@@ Acknowledgments
The research reported in this paper was supported by a Google Focused Research Award on ”Learning to Negotiate Answers in Multi-Pass Semantic Parsing”.
Appendix A Modified Environment
@@
Appendix B Landmark Extraction
The landmarks mentioned in the navigation instructions are extracted before the run starts. We do this by a separate prompt that we feed to GPT-3. The prompt for Map2seq instructions is shown in Figure 8 and the one for Touchdown in Figure 9. It provides five instructions texts paired with a list of extracted landmarks as in-context examples. The example landmark lists were compiled by hand and the same prompt is used for each instance. There are no gold annotations for extracted landmarks and as such no quantitative evaluation is possible. In Figure 10 we show landmarks extracted by GPT-3 using this prompt.
Appendix C Landmark Scorer
We show the CLIP score distribution and panorama views at certain thresholds for additional landmarks in Figure 6.
Appendix D Finished Prompt Sequence
In Figure 7 we show a finished prompt sequence for a given navigation instance. The agent predicted ${\tt{STOP}}$ at timestep 14 and thus finished the trajectory. In the depicted case the agent followed the correct route and successfully completed the navigation objective. For visualization purposes the trajectory is shortened. On average the routes in Touchdown and Map2seq require 40 steps to be completed. This also means the agents has to evaluate 200 panorama views for each navigation instance.

### v2

@@ 6.3 Few-Shot Learning Results
Figure 5: Key point accuracy (KPA) for 2-shot in-context learning of large language models with increasing parameter count. The ${\tt{FORWARD}}$-Only baseline predicts walking forward until the average trajectory length is reached and performs better than predicting random directions.
[…]
@@ 6.4 Finetuning Results
To further explore the capabilities of the proposed LLM agent, we finetune LLaMa-7b on all training instances of the respective dataset, denoted by VELMA-FT in Table 2. Each training instance is the full text sequence that is produced by following the gold path. The visibility of landmarks is determined by the landmark scorer during training because gold annotations are not available. There are 6,770 training instances for Touchdown and 5,737 for Map2seq. We finetune for 20 epochs using LoRA (Hu et al. 2022) to adapt query, key and value projections of the attention layer as well as input and output projection of each transformer layer. The best model is selected by task completion on the development set. The resulting agent outperforms the previous state-of-the-art model ORAR* by 10% and 16% relative TC rate. Comparing ORAR* which fuses image features at the vector level to VELMA-FT which finetunes on verbalizations of observations, shows that the text-based environment observations are less prone to overfitting.
@@
Touchdown Map2seq
Image Model SPD↓ TC↑ SPD↓ TC↑
no image 26.7 ±0.4 15.0 ±0.6 9.1 ±0.2 37.8 ±1.0
CLIP 20.2 ±0.3 20.8 ±0.5 8.8 ±0.5 39.2 ±0.5
OpenCLIP 18.3 ±0.4 23.4 ±0.2 8.7 ±0.0 41.3 ±0.9
A navigation task is successfully completed if the agent stops at either the goal location or an adjacent neighboring node. Training the agent with teacher-forcing to exactly follow the gold route penalizes the agent for stopping one step short or one step past the target node, despite accomplishing the navigation objective. Furthermore, the agent can not learn to recover from incorrect decisions during inference. We thus train the agent to directly optimize the TC metric while also feeding it its own actions during training, called VELMA-RBL in Table 2. The procedure for VELMA-RBL is inspired by response-based learning (Clarke et al. 2010) and imitation learning (Ross, Gordon, and Bagnell 2011) and is outlined in Algorithm 1. The loss for an instance at training step $j$ is either computed by teacher forcing the gold action sequence $\hat{\mathbf{a}}$, or by student forcing, determined by a mixing parameter $\lambda$. In student forcing, the actions decoded by the current model weights $\theta_{j}$ are executed instead of the gold actions. If this trajectory ends within one neighboring node of the target location, the predicted action sequence $\mathbf{a}_{j}$ is considered correct and used as the reference to train the agent. If the agent stops at the wrong location, an oracle path is computed to provide the optimal counterfactual action at each step in the trajectory. In our case, the oracle’s optimal next action is computed as the shortest path to the goal location. We set $\lambda=0.5$ to collect training losses in a batch evenly from both training strategies. Manually inspecting trajectories produced by the trained agent, we found improvements of following instructions that have stopping criteria like ”Stop a few steps before Y.” or ”Stop at X. If you see Y you have gone too far.”. In both cases, the agent learned to walk past the uncertain stopping location and to invoke the ${\tt{TURN\_AROUND}}$ action in order to walk back once landmark Y appeared. The described training procedure leads to a significant increase of task completion rate by 2.9 and 7.5 for Touchdown and Map2seq, respectively. Overall, our contributions in this work amount to a relative increase of task completion by 77% and 57% over the previously reported state-of-the-art for urban VLN on the Touchdown and Map2seq datasets.
@@ 7 Conclusion
We introduced VELMA, an agent for urban vision and language navigation, which utilizes a large language model to infer its next action. The LLM is continuously queried with a text prompt that verbalizes the task description, navigation instructions, visual observations, and past trajectory of the agent. In order to include observed landmarks in the prompt, we propose an unsupervised pipeline that extracts landmarks from the instructions and determines their visibility in the current panorama view based on thresholded CLIP scores. We evaluate the embodied LLM agent in a modified version of the commonly used Touchdown environment based on Street View. One proposed modification is fixing a problem at intersections that led to incorrect alignments of action sequences, and another modification adds the ${\tt{TURN\_AROUND}}$ action which provides a more intuitive way to communicate with the environment. The proposed agent achieves strong few-shot in-context learning results of 10 and 23 task completion rates for Touchdown and Map2seq, respectively, and yields new state-of-the-art results of 26 and 47 task completion rates when finetuned on the full training set. The finetuning results show that verbalization is not an inherent limitation for this task and in-context learning with better base models or improved prompting techniques could outperform our reported few shot results.
@@ Acknowledgments
The research reported in this paper was supported by a Google Focused Research Award.
Appendix A Finetuning Details
In the finetuning experiments, we use the official LLaMA-7b weights and finetune a LoRA adapter for q_proj and v_proj. LoRA hyperparameter are set to $r=8$, $alpha=16$, $dropout=0.05$ and no bias. We use Adam (Kingma and Ba 2015) as the optimizer with a learning rate of $0.0003$, warmup ratio of 0.1 and linear decay. The batch size is 16 and we train for 20 epochs. We use greedy decoding for all experiments.
Appendix B Modified Environment
@@
Appendix C Landmark Extraction
The landmarks mentioned in the navigation instructions are extracted before the run starts. We do this by a separate prompt that we feed to GPT-3. The prompt for Map2seq instructions is shown in Figure 8 and the one for Touchdown in Figure 9. It provides five instructions texts paired with a list of extracted landmarks as in-context examples. The lists of example landmarks were compiled by hand and the same prompt is used for each instance. There are no gold annotations for extracted landmarks and as such no quantitative evaluation is possible. In Figure 10 we show landmarks extracted by GPT-3 using this prompt.
Appendix D Landmark Scorer
We show the CLIP score distribution and panorama views at certain thresholds for additional landmarks in Figure 6. Some navigation instructions refer to the flow up traffic when orientating the agent in the beginning of Touchdown instances, e.g. ”Orientate yourself with against the flow of traffic…”. To support this kind instructions, we score the phrases ”the front view of a vehicle” and ”the rear view of a vehicle” once, before the start. Whichever phrase scores higher with the initial perspective, determines if the agent is facing against the traffic or with the flow of traffic respectively. This traffic flow prediction is then provided as an environment observation string before the first step of the agent.
Appendix E Full Prompt Sequence
In Figure 7 we show a full prompt sequence for a given navigation instance. The agent predicted ${\tt{STOP}}$ at timestep 14 and thus finished the trajectory. In the depicted case the agent followed the correct route and successfully completed the navigation objective. For visualization purposes the trajectory is shortened. On average the routes in Touchdown and Map2seq require 40 steps to be completed. This also means the agents has to evaluate 200 panorama views for each navigation instance.